# Opportunities and Challenges of Generative-AI in Finance

### v1

@@ Abstract
Machine Learning and data mining have created widespread impact across various domains. However, these techniques are limited in their ability to reason, understand and generalize w.r.t language specific tasks. The aforementioned challenges were overcome, with the advancement of LLM’s/Gen-AI. Gen-AI techniques are able to improve understanding of context and nuances in language modeling, translation between languages, handle large volumes of data, provide fast, low-latency responses and can be fine-tuned for various tasks and domains.
In this manuscript, we present a comprehensive overview of the applications of Gen-AI techniques in the finance domain. In particular, we present the opportunities and challenges associated with the usage of Gen-AI techniques in finance. We also illustrate the various methodologies which can be used to train Gen-AI and present the various application areas of Gen-AI techniques in the finance ecosystem.
@@ I Introduction
Machine Learning and data mining have created widespread impact across several domains such as healthcare, payments, Internet of Things (IoT), automation and e-commerce [96]. The success of these techniques can be attributed to their ability to identify and generalize patterns across large amounts of data. In particular, these techniques have led to automation of repetitive tasks, improved precision and accuracy, enhanced decision making, personalized and tailored customer experience and operational scalability across various methodologies. Clustering [52], future prediction [56], anomaly detection [57], time-series monitoring [58] and graph-based learning are some examples of data mining and machine methodologies widely used across domains. However, machine learning and data mining techniques such as Support Vector Machine’s (SVMs) [53], gradient boosted trees [55] and logistic-regression [54] based techniques are limited in their ability to reason, understand and generalize w.r.t. language specific tasks.
@@
The aforementioned challenges were overcome, with the advancement of Large Language Models (LLM’s). LLM’s are able to improve understanding of context and nuances in language modeling [98], translation between languages, handle large volumes of data, provide fast and low-latency responses and can be fine-tuned for various tasks and domains. The advancement was sparked by the development of transformer architecture, which is based solely on attention mechanisms, dispensing with recurrence and convolutions [47]. Experiments have been performed to demonstrate how LLM’s have superior quality while requiring significantly less time to train [99, 100, 101, 102]. Please note that we will use the words Gen-AI and LLM’s interchangeably from here on.
@@
There are several opportunities associated with the usage of Generative Artificial Intelligence (Gen-AI) based technologies in finance[12]. Broadly speaking, they can be used across multiple avenues such as interactive (e.g. chatbots), assistive (e.g. easing payment related activities, right card to use), educative (e.g. enabling users to understand finance concepts) and advisory (e.g. trading assistant).
@@
Along with the opportunities, there lies several challenges associated with the widespread adoption of Gen-AI in finance. Some of these challenges are originating from the scarcity of private and sensitive quality data available for training, issues stemming from pre-training and fine-tuning of financial LLM’s, inference latency, computational costs associated with deploying these models in production, exorbitant cost associated with the pricing of API’s available for commercial usage, inherent biases and privacy embedded within these models [88, 103].
@@
There are several avenues via which Gen-AI models can be trained and developed for financial use-case. Users can leverage either open source models or external service providers for their tasks [104]. Alternatively, one can explore the usage via prompt engineering i.e. zero-shot and few-shot learning. In scenarios, when utilizing Gen-AI techniques out of the box is not working for the task at hand, then one could leverage fine-tuning (e.g. instruction specific, task-specific or parameter efficient) to train in order to perform better at the specific tasks the user has in mind. To overcome issues stemming from hallucination and exponential increase in the number of parameters, several interesting solutions have been proposed and will be discussed later.
@@
Gen-AI has also been widely applied in the finance industry with its application to customer service and support (e.g. sentiment analysis, chat-bots), text summarization and assistance (e.g. summaries from large text, recommending knowledge from large corpora), auto-filling forms, risk management ( e.g. market risk analysis, credit risk monitoring, anomaly detection), investment trading ( e.g. quantitative analysis ) and document processing ( e.g. regulatory compliance) [105] [106].
@@
The organization of this paper is as follows. In Section-2, we discuss the various opportunities associated with the usage of Gen-AI in the finance ecosystem. In Section-3, we present the various challenges associated with the usage of these techniques. Further, in Section-4, we illustrate the various methodologies which can be used to train Gen-AI for widespread adoption. Furthermore, in Section-5, we present the various application areas of Gen-AI in finance. Lastly, in Section-6, we conclude this manuscript.
@@ II Opportunities of Gen-AI in finance
In this section we will outline the opportunities associated with the usage of Gen-AI in the finance domain. Broadly, the opportunities could be classified into four major categories i.e. interactive, assistive, educative and advisory.
@@ II-A Interactive
Gen-AI is the state-of-the-art technology for both Task-Oriented Dialog (TOD) [123] and Open-Domain Dialog (ODD), thereby making it the go-to technology for building conversational and interactive applications. McKinsey [68] estimates that Gen-AI could further reduce the volume of human-serviced contacts by up to 50 percent, depending on a company’s existing level of automation. Applying generative AI to customer care functions could increase productivity at a value ranging from 30 to 45% of current function costs. Efficient customer service is one of the competitive advantages finance companies have and often build closed source proprietary Gen-AI systems to gain competitive advantage. Klarna is one of the Buy Now Pay Later (BNPL) companies that has publicly stated that its new customer service tool can do the work equivalent of 700 full time human agents and handles more than 66% of its customer conversations. It has saved the company over \$40M and reduced the ticket resolution times from 11 minutes to 2 minutes [119].
Gen-AI is enabling the paradigm shift towards end-to-end training of conversational agents by eliminating the need for separate components for natural language understanding, dialogue state tracking, and natural language generation. Their inherent ability to ask follow up questions and generate concise explanations of complex financial concepts for any user’s level of understanding is a big value add. These conversational agents can effectively track references to previous utterances, entities, and user preferences, enabling them to generate more coherent, consistent, and personalized responses that are available 24/7 [37]. These are a vast improvement over rule based chat-bots which could respond to a fixed set of queries with a pre-configured set of responses. Fine-tuning and Retrieval Augmented Generation (RAG) [38] can further help LLM’s become more domain specific and/or access more up to date/ relevant information to support customers ranging from routine to complex inquiries while creating more personalized and engaging experiences [39].
Customer agent avatars created dynamically for digital communications can be tailored to match the customer’s preferences, creating a sense of familiarity and trust. Gen-AI could also be used to create appealing backgrounds for debit, credit cards, coupons and gift cards. Many models including Parti [40] can generate high-fidelity photo-realistic images and supports content-rich synthesis involving complex compositions and world knowledge. One could go one step further and create dynamic visually pleasing themes for apps that are more engaging and context specific.
@@ II-B Assistive
GenAI’s ability to assist through digital automation, co-piloting, autofilling, shopping would greatly enable individuals by being a sophisticated assistant. Further, this may greatly assist individuals with mobility and dexterity challenges.
Gen-AI can be trained to call API’s [60] enabling them to execute tasks, automate routines assist people improving the quality of life of individuals.This can help facilitate payments and shopping [42]. A payments co-pilot can help identify opportunities for better payment options between cards with better benefits and suggest relevant offers for the user’s context. This co-pilot can also help choose between different instrument modes like cards, bank accounts and buy now pay later payment modes that the user could qualify for and will help users make better payment choices based on the metadata of the transaction.
Auto-filling forms is another use case that Gen-AI can potentially excel at. Auto-fill is a hard problem due to the dynamic and changing nature of forms across the internet. Writing automation logic to auto-fill forms doesn’t scale and fails soon. Ability of Gen-AI to understand any page content and context via parsing document object model and the user intent provides new opportunities to fill out forms more accurately.
Agents can be built that help to shop by being the digital shopping assistant. LLaSA is one such LLM designed specifically for this purpose [107]. Users could simply describe what they are looking for and the LLM’s could parse the inventory and provide recommendations that closely match the user’s needs. These agents could be given personalities that make the brand unique and an animated digital persona for the assistant could also be generated using many of the available text to avatar animation models [41].
@@ II-C Educative
Gen-AI forms the foundation around which tools that help improve financial literacy can be built. This would lead to creation of applications that can guide, educate and empower users.
While the use of Machine Learning (ML) and Artificial Intelligence (AI) for estimating and recommending optimal investment strategies are not new [4, 13], Gen-AI’s ability to ingest vast amounts of financial data to analyze market sentiment, news and other data sources can help inform better trading strategies [3, 2].
Gen-AI can be used to perform deep dives into existing financial data about a customer or company, which can have variety of applications for a company or individual, ranging from uncovering financial trends that could lead to cost efficiency measures, investing opportunities and at a minimum, provide a quick access knowledge base for financial advisors.
Chatbots [43] extensively trained on the finance data helps users understand complex topics like mortgage planning, investments like stock options, mutual funds and retirement planning. Their ability to have multi turn conversations enable them to answer questions to all kinds of users i.e. from beginners to advanced.
@@ II-D Advisory
Financial institutions leverage the capabilities of Gen-AI to build powerful tools for financial advisory applications (e.g. risk assessment, pro-active detection, summarization and recommendation).
Gen-AI can help to build sophisticated risk models [44] by processing vast amounts of financial data to identify anomalous patterns of fraud. Such models can be used for better credit scoring, risk assessment for loans, to identify fraudulent activities in payment systems and to forecast financial risks using chain of thought prompting[46] to elicit better reasoning and accuracy.
Further, It can be used to create customized security alerts based on transaction history and behavior which could help reduce the risk of fraud. SMS based scams [smishing] are on the rise, In 2021 alone, SMS phishing has caused a huge \$44 billion in losses just within the United States [48]. Unfortunately, generative AI chat-bots are often abused to create such smishing campaigns [49]. Ironically, AI agents themselves can provide a path forward to tackle these issues by parsing, analyzing and flagging such messages. Such features could be built into the operating system of the phone helping to proactively detect and alert the users of a potential scam.
LLM’s showcase a remarkable capacity for document summarization. By leveraging extensive pre-training on diverse text corpora, they can generate concise yet informative summaries by extracting salient and relevant information about the topic that a user is interested in. They can identify key themes, track discourse structure and distill the core message all while capturing the essence of the original text. Further, this capability can be used for text summarization tasks and explanation of signup forms, legal agreements, financial terms and conditions.
@@ III Challenges of using Gen-AI in finance
Gen-AI has the potential to be a transformational technology in finance and payments. However, this area is not without its unique challenges. We have categorized these challenges into the following categories[92].
III-A Data Availability
There are two challenges when it comes to data used to train Gen-AI models.
The first challenge stems from an abundance of publicly available data along with simultaneous scarcity of private and sensitive quality data [16, 14, 12]. Finance and payments industry has very stringent reporting regulations and as a result a lot of data is available publicly from various financial institutions. However, when it comes to the internal and private data of financial institutions, there is a scarcity of quality data. When Gen-AI models are trained on this abundant public and scarce private data, it is not clear which data gets used for specific decision making. For example, auto insurance risk assessment for an individual could be made based on publicly available data about the individual, whereabouts of their vehicle while the individual might not be even aware that this data is available to their insurance company [10].
Secondly, even when the data quality bar is met, many attributes of the financial data are considered privacy sensitive and are subject to stringent regulations on its use [14].
The above two challenges can be partially overcome by the use of synthetic data. Synthetic data[109, 110] is one of the solutions financial institutions are using to solve the problem of privacy and quality of data. Synthetic data refers to data that is artificially generated for the training purposes and very representative of the real world data and yet is not really from a real world entity [16]. While such data might help institutions mitigate the data privacy issue it also may deprive them of the use of emergent properties of Gen-AI that can uncover hidden patterns and trends in the data[111].
@@ III-B Challenges in fine tuning
The importance of fine tuning Gen-AI models is well understood [18]. Gen-AI models require extensive fine tuning when they are applied to specific domains such as finance. Fine tuning is often performed using raw data where input and expected output is both fed to the model, another approach is Instruct fine-tuning [19] where the expected output is not provided directly but rather more explicit instructions are provided to the model.
FinBERT [17], a LLM for finance domain is a good data point that highlights the importance of pre-training and fine tuning. FinBERT’s insight was that financial vocabulary is an important factor in accuracy of financial models. Such vocabulary needs to be carefully embedded into the model of the pre-training and not just fine-tuning (Example, associating “bank” with “lending” more than “river”).
FinGPT which uses LoRA (Low Rank Adaptation) technique [6] to finetune available general purpose open source LLM’s [97] (e.g. LLaMA) and apply them to finance domains shows the limitations and strengths of fine tuning. Research [20] has reported that fine tuned models based on open source models like LLaMA are able to outperform BloombergGPT on finance related tasks even with a modest amount of training data of 50K samples[20]. However such models were fine tuned for very specific types of financial tasks.
@@
Current research suggests[20] that availability of high quality human labeled data is critical for fine tuning. In the finance domain producing human labeled data is harder due to challenges around regulations, privacy and availability of humans with such specialized knowledge. Additionally research [20] predicts that the cost of fine tuning an existing open source model could be more than \$30,000.
III-C Latency and time validity of results
Latency refers to the amount of time a system takes to produce output once you provide it an input. In the financial domain, latency is of high importance because the more transactions happen per second, more revenue is generated. In some of the cutting edge high frequency trading platforms, lowering latencies by 1 millisecond can lead to additional revenue around \$100M [21].
Latency is also a concern around adoption of Gen-AI technologies. Gen-AI models could take a significant amount of time to produce their output. Benchmarking of various Gen-AI models shows that the inference latency varies a lot based on the context window[21]. In finance applications context windows are expected to be much longer especially when the context might include past data specific to the query, time series data and various signals required for inference. Hence any Gen-AI based synchronous operation will end up impacting the latency significantly [114].
Another aspect that must be considered in conjunction with latency is the time validity of the data. In financial applications where Gen-AI is expected to work off the real time data, the real time data might have very limited time validity. For example stock prices change very rapidly and might be invalid by the time the Gen-AI model is inferred.
III-D Computation costs
Computational costs associated with training/inferring Gen-AI models are a source of another challenge III-B towards widespread adoption. Estimating aforementioned costs is an evolving field with rapidly changing baselines [112].
Credit card payments hover around 2B transactions per day. Stock market transactions and other banking transactions are even more [22]. Such transactions are enabled by different types of institutions, which incur a cost every time a transaction executes. Introducing Gen-AI based computations would further increase cost.
The pricing models of Gen-AI APIs available for commercial use vary from \$5 to \$15 per million input tokens. If we assume that each credit card transaction might involve 500 tokens, this then comes to around the cost of \$0.0025 per transaction or \$5M for 2B transactions per day in additional expenses [87].
If we apply similar math to stock market predictions then the token window is going to be a lot longer and costs even more. Hence, managing computational costs is going to be one of the most critical challenges to use Gen-AI in finance in future. As of today, no major player has published any real world data around economics of using Gen-AI in finance.
III-E Risk considerations
Gen-AI systems have some well known data privacy and bias issues. In finance, where regulations and laws govern nearly all activities, these problems become extremely critical to manage [88].
Gen-AI models are trained on massive amounts of data, some of which could be the user’s private data. Even when anonymized some of this data could be linked to specific users by clever prompt engineering and inference. This data leakage is one of the biggest risks of using Gen-AI in finance [115].
Certain data used in training might later be discarded by the entity that owns it or the entity might request it to be deleted[95]. However, when such data is used in training; even though it is not explicitly stored in the models, it is still remembered by the model. This raises privacy and regulatory concerns[88]. In the era of right to be forgotten [116] regulations, this is a major risk.
Besides data leakage, embedded bias in finance is another major problem. Different regions have different laws governing what kind of information can or can not be used for making decisions about say creditworthiness of an individual[89]. Past research [90] has shown that pricing of say housing market is affected by human bias around race. Gen-AI models trained on such data might perpetuate such biases without even realizing about their existence.
The challenges around data privacy and ethics go hand in hand with regulatory compliance of FIs. As of today, we do not have comprehensive regulations [91] on use of Gen-ai in finance but we can expect such regulations to emerge soon. It is important that companies design their Gen-AI strategies keeping these challenges in mind.
Gen-AI’s domain independent problems such as hallucination and inconsistent reasoning[94] also is a problematic aspect in finance where wrong information can lead to significant financial losses. Various LLM’s for example have been found to be highly inconsistent when it comes to banking related information[93].
IV Gen-AI Methodologies In Finance Ecosystem

### v3

@@ Abstract
Gen-AI techniques are able to improve understanding of context and nuances in language modeling, translation between languages, handle large volumes of data, provide fast, low-latency responses and can be fine-tuned for various tasks and domains.
In this manuscript, we present a comprehensive overview of the applications of Gen-AI techniques in the finance domain. In particular, we present the opportunities and challenges associated with the usage of Gen-AI techniques. We also illustrate the various methodologies which can be used to train Gen-AI techniques and present the various application areas of Gen-AI technologies in the finance ecosystem.
@@ I Introduction
Financial sector has historically applied traditional machine learning technologies to address domain specific challenges. For example, clustering [1], predictive analysis [2], anomaly detection [3], time-series monitoring [4], and graph-based learning have found extensive use for finance-specific problems such as detecting trends, prediction, risk modeling, and better data representation. However, Generative AI (Gen-AI) technologies, such as large language models (LLMs), have a superior ability to understand tasks that involve language compared to any traditional techniques. Kashyap et al. [5] discuss that LLMs can effectively perform traditional machine learning tasks and, when used correctly, can even outperform traditional methods such as classification.
@@
Gen-AI techniques are able to improve understanding of context and nuances in language modeling [6], translation between languages, handle large volumes of data, provide fast and low-latency responses and can be fine-tuned [7] for various tasks and domains. The advancement was sparked by the development of transformer architecture, which is based solely on attention mechanisms, dispensing with recurrence and convolutions [8]. Experiments have been performed to demonstrate how Gen-AI techniques have superior quality while requiring significantly less time to train [9, 10].
@@
There are several opportunities associated with the usage of Gen-AI technologies in finance [11]. Broadly speaking, they can be used across multiple avenues such as interactive (e.g. chatbots), assistive (e.g. easing payment related activities, right card to use), educative (e.g. enabling users to understand finance concepts) and advisory (e.g. trading assistant).
@@
Along with the opportunities, there lies several challenges associated with the widespread adoption. Some of these challenges originate from the scarcity of private and sensitive quality data available for training, issues stemming from pre-training and fine-tuning, inference latency, computational costs associated with deploying these models in production, exorbitant cost associated with the pricing of APIs available for commercial usage and the inherent biases embedded within these models [12, 13].
@@
There are several avenues via which Gen-AI technologies can be trained and developed for financial use-case. Users can leverage either open source models or external service providers for their tasks [14]. Alternatively, one can explore the usage via prompt engineering i.e. zero-shot and few-shot learning. In scenarios, when utilizing Gen-AI techniques out of the box is not working for the task at hand, then one could leverage fine-tuning (e.g. instruction specific, task-specific or parameter efficient) to train in order to perform better at the specific tasks the user has in mind.
@@
Gen-AI techniques also been widely applied in the finance industry with its application to customer service and support (e.g. sentiment analysis, chat-bots), text summarization and assistance (e.g. summaries from large text, recommending knowledge from large corpora), auto-filling forms, risk management (e.g. market risk analysis, credit risk monitoring, anomaly detection), investment trading (e.g. quantitative analysis) and document processing (e.g. regulatory compliance) [15] [16].
@@
The organization of this paper is as follows. In Section II, we discuss the various opportunities associated with the usage of Gen-AI in the finance ecosystem. In Section III, we present the various challenges associated with the usage of these techniques. Further, in Section IV, we illustrate the various methodologies which can be used to train Gen-AI for widespread adoption. Furthermore, in Section V, we present the various application areas of Gen-AI in finance. Lastly, in Section VI, we conclude this manuscript.
@@ II Opportunities of Gen-AI in finance
In this section, we outline the opportunities associated with the usage of Gen-AI techniques in the finance domain. Broadly, we classify the opportunities into four major categories i.e. interactive, assistive, educative and advisory.
@@ II-A Interactive
Gen-AI technology has the potential to become a state-of-the-art technology for both Task-Oriented Dialog [17] and Open-Domain Dialog, thereby making it the go-to technology for building conversational and interactive applications. The associated inherent ability with these techniques to ask follow up questions, track references to previous entities, and user preferences makes them more coherent and consistent. They are a vast improvement over rule-based chat-bots which could respond to a fixed set of queries with a pre-configured set of responses. Fine-tuning and Retrieval Augmented Generation (RAG) [18] can further help Gen-AI techniques, become more domain specific and/or access more up to date relevant information to support customers ranging from routine to complex inquiries while creating more personalized and engaging experiences. McKinsey [19] estimates that Gen-AI could further reduce the volume of human-serviced contacts by up to 50 percent, depending on a company’s existing level of automation. Further, Klarna, a Buy Now Pay Later (BNPL) company has observed that using Gen-AI techniques within its customer service tool can lead to a \$40M impact [20].
Gen-AI techniques could also be used to create appealing backgrounds for debit, credit cards, coupons and gift cards. Many models including Parti [21] can generate high-fidelity photo-realistic images and supports content-rich synthesis involving complex compositions and world knowledge. Further, Gen-AI techniques can be used to create dynamic visually pleasing themes for apps that are more engaging and context specific.
@@ II-B Assistive
Gen-AI techniques have the ability to assist through digital automation, co-piloting, auto-filling and shopping.
In particular, agents using Gen-AI techniques, can be trained to execute APIs [22] thereby enabling them to execute tasks, automate routines thereby substantially improving the quality of life with humans having mobility and dexterity challenges. Further, a payments co-pilot can help identify opportunities for better payment options among cards, instrument modes (e.g. cards, bank accounts).
Gen-AI techniques are able to parse and understand page content and hence can help towards auto-filling, which is a complex problem due to the dynamic and changing nature of forms across the internet. Further, agents using Gen-AI technology can be used as digital assistants. LLaSA [23], a digital assistant is able to parse the product inventory and provide recommendations that closely match the user’s needs.
@@ II-C Educative
Gen-AI techniques have the potential to create applications which can guide, educate and empower users w.r.t. financial literacy.
Gen-AI technology can be used to perform deep dives into existing financial data about a customer or company, thereby uncovering financial trends that could lead to cost efficiency measures, or investing opportunities. In particular, Gen-AI techniques have the ability to ingest vast amounts of financial data to analyze market sentiment, news and other data sources to help inform/create better trading strategies [24].
Further, chatbots [25] built using Gen-AI techniques can understand complex topics like mortgage planning and investments strategies (e.g. stock options, mutual funds).
@@ II-D Advisory
Financial institutions can leverage the capabilities of Gen-AI techniques to build powerful tools for financial advisory applications (e.g. risk assessment, pro-active detection, summarization and recommendation).
Gen-AI techniques can be used to build sophisticated machine learning models by processing vast amounts of financial data to identify anomalous patterns of fraud. Such models can be used for better credit scoring, risk assessment for loans, to identify fraudulent activities in payment systems and to forecast financial risks using chain of thought prompting [26] to elicit better reasoning and accuracy. Further, Gen-AI techniques have demonstrated a remarkable ability towards document summarization (i.e. concise yet informative) and identifying key themes.
@@ III Challenges of using Gen-AI in finance
Strict regulations and constraints within the financial domain hinders the development of Gen-AI applications. Further, Gen-AI’s existing limitations present challenges for effective use in financial applications. In this section, we explore these challenges in more detail.
III-A Data
III-A1 Data Constraints
There is a dearth of publicly available financial datasets owing to the sensitive nature of financial data. This limits the availability of open financial Gen-AI models [27]. BloombergGPT, a proprietary Gen-AI model, is the only major model trained for financial tasks from ground up. BloombergGPT is trained on 363 billion tokens of Bloomberg’s own proprietary data and 345 billion tokens of publicly available data [28].
While many financial institutions have access to proprietary data, they still face data availability challenges due to the institutional inertia and privacy related concerns. Kruse et al. [29] conducted a survey of financial service experts to identify key challenges of using Gen-AI and two thirds of the experts indicated the lack of quality training data as the primary obstacle. Further, the same study [29] identified financial institutions’ unwillingness to move the data to cloud as one of the major limiting factors in using Gen-AI in finance. Financial institutions see their data as their core-business and have concerns in moving this data to distributed data centers (or cloud) which is necessary for training.
The lack of quality training data and privacy concerns can be addressed using synthetic data [30]. Synthetic data is artificially generated data that mimics the statistical properties of the real world data. Since the data is artificially generated, it also helps alleviate the concerns around data privacy. Samuel et al. [30] provides a detailed overview of the current state of art of synthetic data generation in finance. However, they also highlight the challenges with synthetic data such as over-fitting and lack of open benchmarks.
III-A2 Nature of Financial Data
Ljung et al. [31] showed that financial data is different from data in other domains because it does not follow Gaussian distribution and hence can not be normalized using state-of-the-art approaches for downstream modeling tasks. This requires the creation of new pre-processing methods for machine learning use-cases. Further, Samuel et al. [30] found that multi-modality and heterogeneity of financial data along with its intricate inter-dependencies are harder to be captured and modeled by Gen-AI technologies.
@@ III-B Challenges in fine tuning
Gen-AI models require substantial fine tuning efforts when they are applied to specific domains such as finance [32]. Fine-tuning is a technique to adapt a pre-trained model for a more specific purpose (e.g. financial applications). Li et al. [33] provided a comprehensive comparison of large language models (e.g. LLaMa) for financial applications and observed that fine tuned models can outperform generic models w.r.t financial tasks.
Further, FinBERT [34], which uses LoRA (Low Rank Adaptation) technique [35] to fine-tune open source LLMs, highlights the importance of pre-training along with fine tuning. In particular, they emphasized that financial vocabulary needs to be carefully embedded into the pre-training stages along with fine-tuning (e.g., associating “bank” with the word “lending” as compared to the word “river”).
@@
Li et al. [33] also highlights the critical nature of high quality human labeled data for fine tuning. They mentioned that, in the finance domain, cost of producing human labeled data is harder and expensive due to challenges around regulations, privacy and availability of humans with specialized knowledge.
III-C Computation costs
The two major contributors to computational costs are training cost and inference cost. Computational costs associated with training Gen-AI models from the ground up can easily run into billions of dollars [36]. However, Xia et al. [36] noted that fine tuning III-B can help reduce this cost for certain applications.
Inference cost associated with the integration of Gen-AI workflows will be high considering the billions of stock market trades, credit card payment and banking transactions within a day [37]. Further, pricing models of Gen-AI APIs available for commercial use vary from \$5 to \$15 per million input tokens. For example, if we assume that each credit card transaction involves 500 tokens, this translates to a cost of \$0.0025 per transaction or \$5 million for 2 billion transactions per day in additional expenses. Bryce et al. [38] provided an overview about how these pricing costs would impact Gen-AI deployment in finance.
III-D Secondary Considerations
Additional considerations of using Gen-AI techniques in finance domain can arise from domain independent issues, embedded domain bias, privacy and regulatory concerns.
Domain independent issues of Gen-AI techniques such as hallucination and inconsistent reasoning [39] using incorrect information might lead to significant financial loss [40].
Embedded domain bias stems from the fact that different regions have different laws governing the kind of information, which can be used for making financial decisions (e.g. credit worthiness) [41]. As a result, Gen-AI techniques trained on such data might further perpetuate biases without being fully aware of their existence.
Data leaks due to sophisticated prompt engineering techniques [42] raises substantial privacy and regulatory concerns [12, 43] surrounding users private data. This complexity further increases due to new privacy regulations (e.g. right to be forgotten) adopted in many regions of Europe [44]. This increase in complexity can be attributed to the fact that a model might still preserve some aspects of the data as part of its learning, even though the underlying data is requested to be deleted.
IV Gen-AI Methodologies